b) Explain possible actions that might be taken to reduce the length of that cycle, 

and the possible disadvantages of each.

c) Assume that the company is negotiating to do business with a new supplier 

who has offered credit terms of 3/15, net 30. The financial manager is planning 

to delay payment for an additional 10 days i.e. to only settle the account after 

40 days. The current bank overdraft rate for the firm is 25% per annum.

Calculate the effective cost of finance provided by this supplier and comment 

on the financial manager’s plans. [3 Marks]

d) An aspect of working capital policy which requires managerial attention is the 

manner in which the items are financed. Discuss aggressive policy in this 

regard.


1
Expert's answer
2021-06-14T16:21:03-0400

b) Adequate cash flow is critical to the success of any business. It is a critical indicator of a company's health. The amount of cash a company has on hand at any given time has an impact on both its day-to-day and long-term procedures and operations. Poor cash flow makes daily commercial activities and operations more difficult for the company, and it also has a negative impact on the time it takes to pay creditors. Reduced cycle duration can also help the company's bottom line by saving money on interest.


c)



= 3% / (100% - 3% ) x 360 / (30 - 15) = 74.23 percent

This contractor's active budget for investment delivery = 74.23 percent. The Executive's Plan: He or she will postpone the payment for another ten days. In such case, the final due date will be = 30 + 10 = 40 days.

The penalty for not taking the discount is then = = 3% / (100 percent - 3% ) x 360 / (40 - 15) = 44.54 percent > 25% = the bank's overdraft rate. As a result, the executive's strategy is flawed; he or she must borrow money and pay the contractor on the 15th day in order to take advantage of the discount.


d) Aggressive policy: Durable assets are funded with temporary coffers. The following is an example: Temporary credits/overdrafts/working capital loans are used to support long-term fixed assets such as machinery and equipment. Inconsistency in resource liability maturity results as a result of this. Moderate: Temporary capital is used to fund short-term assets, whereas long-term capital is used to fund the durable resources.

This working capital management technique is closely monitored. This results in the use of working capital funds, rotational capacity, or an overdraft to fund or bankroll receivables and/or the register.

Conservative: Rather than using long-term funds to support transitory resources, the company borrows on a regular basis to meet its working capital needs, which is inefficient.